How many litres of hydrogen gas are produced at SATP if 7.15 g of magnesium are added to 8.00 g of hydrogen chloride dissolved in water?
1
Expert's answer
2020-06-18T14:45:05-0400

Mg+2HCl--->MgCl2+H2

moles of magnesium taken = 7.15/24 = 0.297 moles

moles of HCl taken = 8/36.5= 0.219 moles

here limiting reagent is HCl so whole HCl will be consumed. so accordingly

moles of H2 produced is half of moles of HCl

moles of H2= (1/2)*0.219 moles= 0.1095 moles


1 mole at STP corresponds 22.4L

so 0.1095 moles of H2 at STP is 22.4*0.1095 L

=2.458L